### v2

@@ V.5.1 Expected rate of DE
As shown in previous studies with XENON1T [61] and LUX [62], the emission rate of delayed electrons are correlated, in time and position, with the previous high-energy events. This part of the background is referred to as a correlated delayed electron background and can be significantly suppressed due to its unique correlation with proceeding high-energy events. We simulate muons passing through the TPC active volume using Geant4, tracking their interaction vertex and the energy deposited along the track. The ionization signals are generated along the muon track, assuming a constant ionization yield of 50 electrons per keV. We then simulated the generation of delayed electrons, assuming that 0.1% of the total ionization electrons from primary interactions are delayed emissions in a time range of 0.1 ms to 2 s, comparable to [61]. The distribution in “delayed time” (time respected to the primary S2) for delayed electrons is assumed to follow a power law with a coefficient of -1.1, as shown in XENON1T [61]. These delayed electrons’ horizontal positions () are assumed to be the same as their primary S2s. Position reconstruction resolution for single electron S2s, which is expected to be 20 mm from optical simulations, is taken into account in this analysis. Ionization S2 signals are grouped together if the time gap between them is below 1 s in order not to split S2 waveforms of events happening at the bottom part of the TPC. As a result, the pile-up of DE (called Pile-up DEs) leads to a significant background in CENS search. In RELICS, we expect more than 10 Hz of muon events in the LXeTPC, leaving a large amount of single electron S2 signals. The expected event rate of a single-electron pile-up background is summarized in Tab. 2. Without implementing appropriate cuts to reduce the rate of these pile-up events, CENS signals with 3 to 6 electrons will be completely overwhelmed.
@@ V.5.3 Pattern classifier
After each S2 signal in the detector, the expected photons to be detected in each channel on the top PMT array are denoted as . is determined by the total area of S2, the relative LCE in that channel, and the PMT responses including detection efficiency. LCE of each PMT is derived with light propagation simulation in the RELICS TPC. It is assumed that the number of photo-electrons received by the PMT channel follows a Poisson distribution with as the mean. The detected photo-electrons are further derived considering the resolution of PMTs to a single photo-electron.
Given the expected and observed light distribution pattern, we calculated the goodness-of-fit likelihood to determine whether the event is from physical interactions or Pile-up DEs. The logarithm of goodness-of-fit likelihood is written as:
where and are expected and observed photo-electrons in a specific channel.
The S2 distribution pattern on top PMT array of the Pile-up DEs rejected by a selection in (figure 12, left) would look like that of physical interactions (figure 12, right). Since both of horizontal position and emission time of DEs are strongly correlated to the previous muon tracks, we further calculate the “Space-time Correlation Coefficient” of each event with previous muons to distinguish Pileup DEs from physical interactions that are independent of any muon events in the RELICS LXeTPC.
@@
We simulated the Pile-up DE events of the detector operating for 8500 hours, and calculated the “Pattern coefficient” and “Spacetime Correlation Coefficient” for each event. The two-dimensional distribution for Pile-up DE and CENS, within an S2 range of 150 to 160 PE, are shown in figure 13.
@@
Figure 14 delineates the score distribution for Pile-up DEs and physical interactions. The final classifier identifies all events with scores exceeding 0 as real physical interactions. In the ROI of S2 ranging from 120 to 240 PE, the acceptance for physical interactions such as CENS is 46.5%. This cut reduced the rate of Pile-up DEs by 4 orders of magnitude in the same ROI, leaving the rate of Pile-up DE background below the expected CENS signal in the whole ROI.
@@ V.6 Summary
The distribution of all the background components is converted to S2 size from recoil energy using RelicsAPT, which is described in Sec. III, and summarized in figure 16. The CENS event rate is overlaid taking into account the signal acceptance. The dominant background contribution comes from the Pile-up DEs, despite its large discrimination power to physical events as described in Sec. V.5. The NR background contribution from CRNs is the second major contribution. The detector materials contribute a relatively low and flat component in ER. The background components’ event rate in 32 kgyr exposure is summarized in Tab. 3.
@@
| Pile-up DE | 1081.9 |
| CRNs | 229.4 |
| Muon-induced neutrons | 1.5 |
| ER | 5.7 |
| Total Background | 1318.4 |
| CENS | 4902.4 |
@@ VI Physics potential
To explore the physics potential of the RELICS detector, this section will discuss its sensitivity to the weak mixing angle and non-standard interactions (NSI) based on the background estimation in section V with 32 kgyr exposure. Several systematic uncertainties are considered, including 5% reactor neutrino flux, 5% ER and NR background uncertainty. The uncertainty of the Qy introduced in Sec. III.1 is the dominant one in our estimation.
@@ VI.1 Weak mixing angle
The of likelihood for measurement is defined as Eq. 7 where index indicates different S2 intervals. and are the observed and expected signal events in each S2 interval, respectively. is the background rate estimation and index runs over background components, including ER and neutrons. The rest are nuisance parameters. and are the multipliers of the ER and neutron backgrounds, respectively. denotes the morphers of charge yields. A Gaussian constraint term is imposed on each nuisance parameter, as summarized in Tab. 4.
The expected constraint on from RELICS is shown in figure 17. The great improvement from the red solid line to the blue dashed line indicates that the charge yield variation is a major contribution to the constraint uncertainty. The sensitivity is expected to be competitive and improve the Dresden-II’s result [63] measured with reactor neutrinos. As illustrated in figure 18, RELICS will provide a measurement on with momentum transfer down to the MeV scale, which is a valuable complementary to the current measurements.
@@ VII Summary
Acknowledgements. We acknowledge CNNC Sanmen Nuclear Power Company for hosting RELICS. We thank Dr. Xun Chen for supporting the BambooMC package and Prof. Kaixuan Ni for fruitful discussions. RELICS is supported by a National Key R&D grant from the Ministry of Science and Technology of China (No. 2021YFA1601600) and a grant from the Natural Science Foundation of China (No. 12275267). Yifei Zhao’s work on RELICS is supported by a grant for undergraduate research from Beijing Natural Science Foundation (No. QY23088).

### v3

@@ V.5.1 Expected rate of DE
As shown in previous studies with XENON1T [71] and LUX [72], the emission rate of delayed electrons are correlated, in time and position, with the previous high-energy events. This part of the background is referred to as a correlated delayed electron background and can be significantly suppressed due to its unique correlation with proceeding high-energy events. We simulate muons passing through the TPC active volume using Geant4, tracking their interaction vertex and the energy deposited along the track. The ionization signals are generated along the muon track, assuming a constant ionization yield of 50 electrons per keV. We then simulated the generation of delayed electrons, assuming that 0.1% of the total ionization electrons from primary interactions are delayed emissions in a time range of 0.1 ms to 2 s, comparable to [71]. The distribution in “delayed time” (time with respect to the primary S2) for delayed electrons is assumed to follow a power law with a coefficient of -1.1, as shown in XENON1T [71]. These delayed electrons’ horizontal positions () are assumed to be the same as their primary S2s. Position reconstruction resolution for single electron S2s, which is expected to be 20 mm from optical simulations, is taken into account in this analysis. Ionization S2 signals are grouped together if the time gap between them is below 1 s in order not to split S2 waveforms of events happening at the bottom part of the TPC. As a result, the pile-up of DE (called Pile-up DEs) leads to a significant background in CENS search. In RELICS, we expect more than 10 Hz of muon events in the LXeTPC, leaving a large amount of single electron S2 signals. The expected event rate of a single-electron pile-up background is summarized in Tab. 2. Without implementing appropriate cuts to reduce the rate of these pile-up events, CENS signals with 3 to 6 electrons will be completely overwhelmed.
@@ V.5.3 Pattern classifier
For an CENS or physical interaction, the expected photo-electrons detected in each channel on the top PMT array are denoted as . is determined by the total area of S2, the relative LCE in that channel, and the PMT responses including detection efficiency. LCE of each PMT is derived with light propagation simulation in the RELICS TPC. It is assumed that the number of photo-electrons received by the PMT channel follows a Poisson distribution with as the mean. The detected photo-electrons are further derived considering the resolution of PMTs to a single photo-electron.
While the “expected” is simulated with multiple electrons generated from the same position, the “observed” photo-electrons in each channel from an “unknown” event is calculated by summing up the simulation from each electron at its corresponding position. Given the expected and observed distribution pattern, we calculated the goodness-of-fit likelihood to determine whether the event is from physical interactions or Pile-up DEs. The logarithm of goodness-of-fit likelihood is written as:
The S2 distribution pattern on top PMT array of the Pile-up DEs rejected by a selection in (figure 12, top) would look like that of physical interactions (figure 12, bottom). Since both the horizontal position and emission time of DEs are strongly correlated to the previous muon tracks, we further calculate the “Space-time Correlation Coefficient” of each event to prior muons to distinguish Pileup DEs from physical interactions that are independent of any muon events in the RELICS LXeTPC.
@@
We simulated the Pile-up DE events of the detector operating for 8500 hours and calculated the “Pattern coefficient” and “Spacetime Correlation Coefficient” for each event. The event distribution in (, ) space slightly varies with different S2 sizes. To clearly illustrate the discrimination power, the Pile-up DE and CENS events in the energy slice of [150, 160] PEs in the (, ) space are shown in figure 13.
@@
Figure 14 delineates the score distribution for Pile-up DEs and physical interactions. The final classifier identifies all events with scores exceeding 0 as real physical interactions. In the ROI of S2 ranging from 120 to 300 PE, the acceptance for physical interactions such as CENS is 46.5%. This cut reduced the rate of Pile-up DEs by 4 orders of magnitude in the same ROI, leaving the rate of Pile-up DE background below the expected CENS signal in the whole ROI.
@@ V.6 Summary
The distribution of all the background components is converted to S2 size in photo-electrons from recoil energy using RelicsAPT, which is described in Sec. III, and summarized in figure 16. The CENS event rate is overlaid taking into account the signal acceptance. The dominant background contribution comes from the Pile-up DEs, despite its large discrimination power to physical events as described in Sec. V.5. The NR background contribution from CRNs is the second major contribution. The detector materials contribute a relatively low and flat component in ER. The background components’ event rate in 32 kgyr exposure is summarized in Tab. 3.
While not utilized for our baseline sensitivity, we are aware that the reactor-OFF period can be used to further constrain our background model [73, 74]. Since the reactor-related background is negligible in the total budget, this data will provide a background-only measurement and dedicated runs can be planned in the future to enhance the signal sensitivity.
@@
| Pile-up DE | 1325.1 |
| CRNs | 339.9 |
| Muon-induced neutrons | 1.7 |
| ER | 21.1 |
| Total Background | 1687.8 |
| CENS | 4639.7 |
@@ VI Physics potential
To explore the physics potential of the RELICS detector, this section will discuss its sensitivity to the weak mixing angle and non-standard interactions (NSI) with neutrino has vector couplings to the up and down quarks based on the background estimation in section V with 32 kgyr exposure. Several systematic uncertainties are considered, including 5% reactor neutrino flux, 10% ER and 5% NR background uncertainty. The uncertainty of the Qy introduced in Sec. III.1 is the dominant one in our estimation.
@@ VI.1 Weak mixing angle
The of likelihood for measurement is defined as Eq. 7 where index indicates different S2 intervals. and are the observed and expected signal events in each S2 interval, respectively. is the background rate estimation and index runs over background components, including ER and neutrons. The rest are nuisance parameters. , and are the multipliers of the ER, neutron and Pile-up DE backgrounds, respectively. denotes the morphers of charge yields. A Gaussian constraint term is imposed on each nuisance parameter, as summarized in Tab. 4.
The expected constraint on from RELICS is shown in figure 17. The great improvement from the red solid line to the blue dashed line indicates that the charge yield variation is a major contribution to the constraint uncertainty. The sensitivity is expected to be competitive and improve the Dresden-II’s result [75] measured with reactor neutrinos. As illustrated in figure 18, RELICS will provide a measurement on with momentum transfer down to the MeV scale, which is a valuable complement to the current measurements.
@@ VII Summary
Acknowledgements. We acknowledge CNNC Sanmen Nuclear Power Company for hosting RELICS. We thank Dr. Xun Chen for supporting the BambooMC package and Prof. Kaixuan Ni for fruitful discussions. RELICS is supported by a National Key R&D grant from the Ministry of Science and Technology of China (No. 2021YFA1601600), a grant from the Natural Science Foundation of China (No. 12275267) and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (No. 22lglj11). Yifei Zhao’s work on RELICS is supported by a grant for undergraduate research from Beijing Natural Science Foundation (No. QY23088).